# Kramer’s Escape Rate and Phase Transition Dynamics in AdS Black Holes with Dark Structures

### v2

@@ Abstract
Keywords: Phase Transition, Free Energy Landscape, Kramer’s Escape Rate, Dark Structures.
1 Introduction
From ancient times to the present, the existence of similar behavioral patterns and recurring elements in various physical systems, ranging from subatomic to cosmic scales, has been abundantly observed. This could be interpreted as a mysterious sign of a fractal structure inherent in the essence and fabric of nature. However, it can also be considered an intentionally hidden shortcut, ingeniously and knowingly embedded in the labyrinth of nature to guide us step by step in deciphering complexities that we cannot grasp all at once. Among the structures that have benefited most from the analogy of behavioral patterns for theoretical modeling, due to their inaccessibility for practical study, are black holes. The comparison of the gravitational field treatment of a black hole and its surrounding world, as a system and environment, with the behavior of a thermodynamic ensemble is indeed one of the most valuable achievements of this modeling. A resemblance that has led to black hole thermodynamics becoming a pioneering and fundamental branch in the evolution of our understanding not only of the structural nature of black holes, but also of the cosmos itself. If we were to highlight the three principal keys that paved the way for the application of thermodynamics and statistical mechanics to the study of black holes, we would undoubtedly mention Hawking temperature, Bekenstein-Hawking entropy, and the conversion of the cosmological constant to pressure. These concepts led to the formulation of the four fundamental laws of black hole thermodynamics [1, 2, 3, 4]. With a slight indulgence, one might also add the exchange and conversion of enthalpy to mass to these golden keys [5].
@@ 1 Introduction
The increasing study of black hole models based on thermodynamics and statistical mechanics has gradually revealed that, apparently, the best model that aligns most closely with current data and computational capabilities, and environmental conditions for studying black hole behavior, is the Van der Waals model (VdW). For instance, in canonical ensembles, a black hole phase transition was found in the background of charged rotating AdS black holes, reminiscent of the liquid-to-gas phase transition of the (VdW) fluid [6, 7, 8]. It was also observed that there is a significant resemblance between the (VdW) liquid-gas system and the charged black hole in the extended phase space [9, 10], or another study showed the critical behavior of charged black hole in fixed charge ensemble coincides exactly with that of (VdW) liquid-gas system [11]. Undoubtedly, phase transition is one of the most fundamental phenomena in classical thermodynamics, which, through extensive study of various structures of matter, has led to a profound understanding of the true behavior of systems in the past. The question that quickly arose was whether, given the (VdW)-like thermodynamic behavior of black holes, these structures, considering their quantum-relativistic structural pattern, could exhibit quasi-material behavior and phase transitions. And if so, what could be the cause of the initiation of this process? The answer to the first part of the question was subsequently obtained, as it was determined that there exists a Hawking-Page phase transition between a large stable black hole and pure thermal radiation [12]. Of course, numerous other studies were conducted that even confirmed the interesting phase transition behaviors for black holes [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. However, the second part of the question remains unclear to us until now.
@@
Although the temperature variation graph of a system can itself be a reliable criterion for phase transition detection, one of the tools that can be used to study the behavioral range of thermodynamic phase transition with greater precision is the free energy landscape. Free energy landscape provides a way to visualize and understand how systems, transition between different states over time. The dependence of free energy on temperature causes temperature changes to significantly alter the free energy landscape. Consequently, studying these apparent changes can provide a better understanding of the system’s behavior and the formation of transitions between different phase states. In thermodynamics, free energy is rewritten in various forms depending on the system’s conditions. However, since the interplay between temperature and pressure is often of primary concern in black hole structures, and since enthalpy offers a better and simpler description of energy transfer at a specific pressure, the Gibbs free energy function is a better candidate among the various definitions of free energy for studying black holes. But the point that was hidden in the study of the phase transition behavior of black holes using Gibbs free energy was that, apparently, in this form of study, no attention was paid to the rate and mode of transition from one state to another. In other words, in these studies, we did not address the following points in the phase transition process from small to large black holes:
• At what rate do these processes occur?
• Is the rate of transition constant across different temperature ranges?
• Does the initial state (small black hole) completely vanish at the end of the transition?
Also, could there be hidden points in the leakage from the small to the large black hole that we are unaware of?
These and similar questions that address the dynamics of phase transitions provide important information about the phase transitions of black holes, which apparently cannot be obtained in the initial analyses of Gibbs free energy. It seems that we were considering the phase transition and its results in a static manner, only at the end points and equilibrium points of Gibbs. Recent studies have observed that phase transition under thermal fluctuations can be examined through the Fokker-Planck equation in non-equilibrium physics, based on the free energy landscape, particularly the Gibbs form. Accordingly, it has been observed that in black holes, a transition occurs from small to large black holes and vice versa, influenced by temperature and the height of the potential barrier in the free energy landscape [29, 30, 31, 32]. According to the studies done, it seemed that the Gibbs free energy is in the form of an effective potential it plays the role of a driving force that leads to the phase transition of the black hole [33]. Based on this, a special focus was formed on the Gibbs energy landscape, which, according to the structure of the Fokker-Planck equation, tried to investigate some kinetics of the phase transition by calculating the mean first passage time in several black holes[34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Of course, other efforts were also made in this direction based on the rewriting of free energy in terms of Landau free energy[57, 58] or thermal potential [59, 60, 61, 62].
@@
One might pose the question: How has the connection between stochastic-form equations and phase transitions, as well as thermodynamic equations, been established? In response, it should be noted that since some thermodynamic processes in the system are driven by stochastic fluctuations, it would be logical to employ the method of analyzing non-equilibrium stochastic processes to study the occurrence of inherently non-equilibrium thermodynamic processes. Building on this foundation, in this paper, we consider two black holes influenced by dark energy, namely:
$\bullet$ Non-Linear Magnetic-Charged AdS black hole surrounded by Quintessence, in the background of Perfect Fluid Dark matter (NLM-C-Q-PFD) [63].
$\bullet$ 4D AdS Einstein-Gauss-Bonnet-Yang-Mills Black hole with a Cloud of Strings (EGB-YM-CS) [64].
@@
Initially, we calculate the various forms of free energy mentioned and demonstrate that all three forms will exhibit nearly identical behavior in the canonical ensemble. Subsequently, using Kramer’s escape rate method, we will examine the characteristics of their first-order phase transition rate in terms of variations in radius, pressure, or temperature. We will then compare our results with those of black holes studied using the same method [59, 60, 61, 62] to determine the influence of the added parameters.
The present paper is organized as follows: In Section 2, we explore the theoretical preliminaries to thoroughly study and introduce various concepts, such as Kramer’s escape rate and different forms of the free energy landscape. This includes the Gibbs free energy landscape, Landau free energy, and thermal potential. In Section 3, we conduct a behavioral comparison of the free energy landscape and Kramer’s escape rate for (NLM-C-Q-PFD) AdS black hole. Section 4 presents an accurate calculation for the (NLM-C-Q-PFD) AdS black hole. In Section 5, we fully analyze the concepts discussed in the previous two sections. Finally, in Section 6, we comprehensively present the results of our work.
2 Theoretical preliminaris
2.1 Kramer’s escape rate
@@
The study of Brownian particle leakage or escape under thermal fluctuations from a stable or metastable potential, known as the Kramers problem, is a fascinating topic that has garnered attention across most branches of physics [62]. This subject, in addition to the diverse results it has yielded throughout its study, has led to the calculation of a relationship known as the Kramers escape rate. We will briefly address this for a better understanding here.
@@
As we can see in Fig (1), the point $x_{min}$ is a local minimum such that particles are in equilibrium at this point and maintain their original position under small perturbations. Without loss of generality and for simplicity, we can point out a few things here. For instance, we assume only one direction of escape, which may occur through the nearest local maximum located at $x_{max}$. On the left side of $x_{min}$, the potential is such that it does not have any additional extremes. After the local maximum at $x_{max}$, on the right side, there may be another local or global minimum, or even the potential may be unbounded below, which is irrelevant. This is because the escape rate should be independent of the shape of the potential on the right side of $x_{max}$. When we add the effect of thermal baths to the system, the particles are affected by the two competing effects of fluctuation and dissipation. Random forces arising from thermal fluctuations push the particle away from its initial position and allow it to eventually climb the potential barrier. On the other hand, damping tends to slow the particle down and makes it more difficult to return to the equilibrium point. There are three possible situations here.
• Very Low Damping: In this scenario, particles experience minimal frictional forces. The escape rate is determined by the energy diffusion regime, where the particle’s inertia plays a significant role. The particle can oscillate many times near the potential well’s boundary before escaping, and the escape rate is relatively low due to the lack of significant damping forces.
• Intermediate Damping: The escape rate in this regime is more complex to describe, as it involves a balance between inertial effects and damping forces. The system oscillates less than before, but it can still exist, and the escape rate is still not significant.
• Over-damped: In the over-damped case, the damping forces are strong, and the system experiences high frictional forces, which dominate over inertial effects. The particle moves slowly away from the potential well, and the escape rate is governed by the rate at which the system can overcome the barrier height compared to the thermal energy. The escape process is akin to a slow, diffusive motion over the barrier, and the escape rate is higher than in the very low or Intermediate damping case and is significant.
Since we intend to simulate this model according to the phase transition behavior of the black hole, then among the above options, the over-damping mode seems to be the appropriate option, since due to the strengthening of the combined effect in this mode, the system is no longer stable and the probability of the particle exiting the well will start to grow. That is, after a sufficient period of time, we can expect the particle to have almost passed through the barrier. The escape rate is closely related to the inverse of the average time required for the first passage of the local maximum potential, which is called the ”mean first passage time”, which was noted in articles [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. It should be noted that the escape rate is different from the diffusion rate to the next minimum [65, 66, 67, 68, 69, 70].
Let us consider the Fokker-Planck equation [67, 70, 71],
@@
2.2 Different shapes for free energy landscape
Free energy landscape provides a way to visualize and understand how systems, transition between different states over time. In simple terms, the free energy landscape is a graphical representation where each point on the landscape corresponds to a possible state of the system. From the vantage point of free energy, the exploration of phase transition conditions becomes accessible through the scrutiny of the emergence and stability of distinct phases. Furthermore, the characterization and analysis of the free energy landscape are profoundly influenced by its relationship with the order parameters. Taking the radius of a black hole’s event horizon as an exemplary order parameter, one can articulate the free energy landscape. Consequently, the phase transition of a black hole can be interpreted as a stochastic thermal fluctuation of said order parameter, offering a novel perspective on the underlying physics [33]. In the following, since we want to study the phase transition from the perspective of the free energy landscape and in a different structure under specific thermal fluctuations, we will introduce different forms of this energy that can help us in this way.
2.2.1 Gibbs Free Energy Landscape
@@
The terminology ’on-shell’ and ’off-shell’ delineates the aforementioned temperature discrepancy. ’Off-shell’ encompasses all transitions occurring at temperature ”T”, which, despite conforming to the laws of thermodynamics, do not necessarily constitute solutions to Einstein’s field equations [33]. It is imperative to recognize that solely the extremal points on the $G_{L}$ signify the authentic black hole phases that comply with the Einstein field equations. Furthermore, within this context, the local maxima and minima are indicative of the unstable and locally stable black hole phases, respectively [72].
2.2.2 Landau free energy
The Landau free energy constitutes another showy path for scrutinizing the phase transition dynamics within the system’s free energy landscape. While using this method assumes heightened significance in the context of second-order phase transitions- wherein the bifurcation of the Landau functional’s local minimum epitomizes the thermodynamic system’s second-order phase transition- it is also pertinent to first-order phase transitions. Specifically, the transmutation of the global minimum is emblematic of a first-order phase transition, underscoring the multifaceted nature of the Landau-free energy in characterizing phase behavior. This function can be written as follows [57],
@@ 2.2.2 Landau free energy
2.2.3 Thermal potential
Another form of landscape energy that can be used to investigate phase transition based on thermal fluctuation is the model that was introduced in [59] under the title of thermal potential.
@@ 2.2.3 Thermal potential
3 Non-linear magnetic-charged AdS black hole surrounded
@@
In this article, we went to black holes that somehow interact with the structure of matter and dark energy, so that we can study the effect of the presence of parameters caused by the dark structure on the first-order phase transition. So the metric of (NLM-C-Q-PFD) black holeis as follows [63],
@@
For the main quantities of this model i.e. mass $M$, Hawking temperature $T_{H}$, entropy $S$, pressure $P$ and volume $V$ we will have (note that as long as there is no specific definition of the radius, it will be $r=r_{H}$ in calculations) [63],
@@
3.0.1 Behavioral comparison of landscape free energy
Now according to equations (2), (3), (4) and (7) for Gibbs and Landau free energy function and also the thermal potential of this model, we will have [^5] Note that in the following relationships, due to the number of parameters, the beginning of the equations is shown with a bold symbol only for better clarity.,

### v3

@@ Abstract
Phase Transition, Free Energy Landscape, Kramer’s Escape Rate, Dark Structures
I Introduction
@@
From ancient times to the present, the existence of similar behavioral patterns and recurring elements in various physical systems, ranging from subatomic to cosmic scales, has been abundantly observed. This could be interpreted as a mysterious sign of a fractal structure inherent in the essence and fabric of nature. However, it can also be considered an intentionally hidden shortcut, ingeniously and knowingly embedded in the labyrinth of nature to guide us step by step in deciphering complexities that we cannot grasp all at once. Among the structures that have benefited most from the analogy of behavioral patterns for theoretical modeling, due to their inaccessibility for practical study, are black holes. The comparison of the gravitational field treatment of a black hole and its surrounding world, as a system and environment, with the behavior of a thermodynamic ensemble is indeed one of the most valuable achievements of this modeling. A resemblance that has led to black hole thermodynamics becoming a pioneering and fundamental branch in the evolution of our understanding not only of the structural nature of black holes, but also of the cosmos itself. If we were to highlight the three principal keys that paved the way for the application of thermodynamics and statistical mechanics to the study of black holes, we would undoubtedly mention Hawking temperature, Bekenstein-Hawking entropy, and the conversion of the cosmological constant to pressure. These concepts led to the formulation of the four fundamental laws of black hole thermodynamics 1 ; 2 ; 3 ; 4 . With a slight indulgence, one might also add the exchange and conversion of enthalpy to mass to these golden keys 5 .The increasing study of black hole models based on thermodynamics and statistical mechanics has gradually revealed that, apparently, the best model that aligns most closely with current data and computational capabilities, and environmental conditions for studying black hole behavior, is the Van der Waals model (VdW). For instance, in canonical ensembles, a black hole phase transition was found in the background of charged rotating AdS black holes, reminiscent of the liquid-to-gas phase transition of the (VdW) fluid 6 ; 7 ; 8 . It was also observed that there is a significant resemblance between the (VdW) liquid-gas system and the charged black hole in the extended phase space 9 ; 10 , or another study showed the critical behavior of charged black hole in fixed charge ensemble coincides exactly with that of (VdW) liquid-gas system 11 . Undoubtedly, phase transition is one of the most fundamental phenomena in classical thermodynamics, which, through extensive study of various structures of matter, has led to a profound understanding of the true behavior of systems in the past. The question that quickly arose was whether, given the (VdW)-like thermodynamic behavior of black holes, these structures, considering their quantum-relativistic structural pattern, could exhibit quasi-material behavior and phase transitions. And if so, what could be the cause of the initiation of this process? The answer to the first part of the question was subsequently obtained, as it was determined that there exists a Hawking-Page phase transition between a large stable black hole and pure thermal radiation 12 . Of course, numerous other studies were conducted that even confirmed the interesting phase transition behaviors for black holes a ; b ; c ; d ; e ; f ; g ; h ; i ; j ; k ; m . However, the second part of the question remains unclear to us until now.
@@
Although the temperature variation graph of a system can itself be a reliable criterion for phase transition detection, one of the tools that can be used to study the behavioral range of thermodynamic phase transition with greater precision is the free energy landscape. Free energy landscape provides a way to visualize and understand how systems, transition between different states over time. The dependence of free energy on temperature causes temperature changes to significantly alter the free energy landscape. Consequently, studying these apparent changes can provide a better understanding of the system’s behavior and the formation of transitions between different phase states. In thermodynamics, free energy is rewritten in various forms depending on the system’s conditions. However, since the interplay between temperature and pressure is often of primary concern in black hole structures, and since enthalpy offers a better and simpler description of energy transfer at a specific pressure, the Gibbs free energy function is a better candidate among the various definitions of free energy for studying black holes. But the point that was hidden in the study of the phase transition behavior of black holes using Gibbs free energy was that, apparently, in this form of study, no attention was paid to the rate and mode of transition from one state to another. In other words, in these studies, we did not address the following points in the phase transition process from small to large black holes: At what rate do these processes occur? Is the rate of transition constant across different temperature ranges? Does the initial state (small black hole) completely vanish at the end of the transition? Also, could there be hidden points in the leakage from the small to the large black hole that we are unaware of? These and similar questions that address the dynamics of phase transitions provide important information about the phase transitions of black holes, which apparently cannot be obtained in the initial analyses of Gibbs free energy. It seems that we were considering the phase transition and its results in a static manner, only at the end points and equilibrium points of Gibbs. Recent studies have observed that phase transition under thermal fluctuations can be examined through the Fokker-Planck equation in non-equilibrium physics, based on the free energy landscape, particularly the Gibbs form. Accordingly, it has been observed that in black holes, a transition occurs from small to large black holes and vice versa, influenced by temperature and the height of the potential barrier in the free energy landscape 17 ; 18 ; 19 ; 20 . According to the studies done, it seemed that the Gibbs free energy is in the form of an effective potential it plays the role of a driving force that leads to the phase transition of the black hole 21 . Based on this, a special focus was formed on the Gibbs energy landscape, which, according to the structure of the Fokker-Planck equation, tried to investigate some kinetics of the phase transition by calculating the mean first passage time in several black holes22 ; 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 ; 37 ; 38 ; 39 ; 40 ; 41 ; 42 ; 43 ; 44 . Of course, other efforts were also made in this direction based on the rewriting of free energy in terms of Landau free energy45 ; 46 or thermal potential 47 ; 48 ; 49 ; 50 .
@@
One might pose the question: How has the connection between stochastic-form equations and phase transitions, as well as thermodynamic equations, been established? In response, it should be noted that since some thermodynamic processes in the system are driven by stochastic fluctuations, it would be logical to employ the method of analyzing non-equilibrium stochastic processes to study the occurrence of inherently non-equilibrium thermodynamic processes. Building on this foundation, in this paper, we consider two black holes influenced by dark energy, namely: $\bullet$ Non-Linear Magnetic-Charged AdS black hole surrounded by Quintessence, in the background of Perfect Fluid Dark matter (NLM-C-Q-PFD) 51 . $\bullet$ 4D AdS Einstein-Gauss-Bonnet-Yang-Mills Black hole with a Cloud of Strings (EGB-YM-CS) 52 .
@@
Initially, we calculate the various forms of free energy mentioned and demonstrate that all three forms will exhibit nearly identical behavior in the canonical ensemble. Subsequently, using Kramer’s escape rate method, we will examine the characteristics of their first-order phase transition rate in terms of variations in radius, pressure, or temperature. We will then compare our results with those of black holes studied using the same method 47 ; 48 ; 49 ; 50 to determine the influence of the added parameters. The present paper is organized as follows: In Section 2, we explore the theoretical preliminaries to thoroughly study and introduce various concepts, such as Kramer’s escape rate and different forms of the free energy landscape. This includes the Gibbs free energy landscape, Landau free energy, and thermal potential. In Section 3, we conduct a behavioral comparison of the free energy landscape and Kramer’s escape rate for (NLM-C-Q-PFD) AdS black hole. Section 4 presents an accurate calculation for the (NLM-C-Q-PFD) AdS black hole. In Section 5, we fully analyze the concepts discussed in the previous two sections. Finally, in Section 6, we comprehensively present the results of our work.
II Theoretical preliminaris
II.1 Kramer’s escape rate
@@
The study of Brownian particle leakage or escape under thermal fluctuations from a stable or metastable potential, known as the Kramers problem, is a fascinating topic that has garnered attention across most branches of physics 50 . This subject, in addition to the diverse results it has yielded throughout its study, has led to the calculation of a relationship known as the Kramers escape rate. We will briefly address this for a better understanding here.
@@
As we can see in Fig. (1), the point $x_{min}$ is a local minimum such that particles are in equilibrium at this point and maintain their original position under small perturbations. Without loss of generality and for simplicity, we can point out a few things here. For instance, we assume only one direction of escape, which may occur through the nearest local maximum located at $x_{max}$. On the left side of $x_{min}$, the potential is such that it does not have any additional extremes. After the local maximum at $x_{max}$, on the right side, there may be another local or global minimum, or even the potential may be unbounded below, which is irrelevant. This is because the escape rate should be independent of the shape of the potential on the right side of $x_{max}$. When we add the effect of thermal baths to the system, the particles are affected by the two competing effects of fluctuation and dissipation. Random forces arising from thermal fluctuations push the particle away from its initial position and allow it to eventually climb the potential barrier. On the other hand, damping tends to slow the particle down and makes it more difficult to return to the equilibrium point. There are three possible situations here. Very Low Damping: In this scenario, particles experience minimal frictional forces. The escape rate is determined by the energy diffusion regime, where the particle’s inertia plays a significant role. The particle can oscillate many times near the potential well’s boundary before escaping, and the escape rate is relatively low due to the lack of significant damping forces. Intermediate Damping: The escape rate in this regime is more complex to describe, as it involves a balance between inertial effects and damping forces. The system oscillates less than before, but it can still exist, and the escape rate is still not significant. Over-damped: In the over-damped case, the damping forces are strong, and the system experiences high frictional forces, which dominate over inertial effects. The particle moves slowly away from the potential well, and the escape rate is governed by the rate at which the system can overcome the barrier height compared to the thermal energy. The escape process is akin to a slow, diffusive motion over the barrier, and the escape rate is higher than in the very low or Intermediate damping case and is significant. Since we intend to simulate this model according to the phase transition behavior of the black hole, then among the above options, the over-damping mode seems to be the appropriate option, since due to the strengthening of the combined effect in this mode, the system is no longer stable and the probability of the particle exiting the well will start to grow. That is, after a sufficient period of time, we can expect the particle to have almost passed through the barrier. The escape rate is closely related to the inverse of the average time required for the first passage of the local maximum potential, which is called the ”mean first passage time”, which was noted in articles 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 ; 35 ; 36 ; 37 ; 38 ; 39 ; 40 ; 41 ; 42 ; 43 ; 44 . It should be noted that the escape rate is different from the diffusion rate to the next minimum 53 ; 54 ; 55 ; 56 ; 57 ; 58 . Let us consider the Fokker-Planck equation 55 ; 58 ; 59 ,
@@
II.2 Different shapes for free energy landscape
Free energy landscape provides a way to visualize and understand how systems, transition between different states over time. In simple terms, the free energy landscape is a graphical representation where each point on the landscape corresponds to a possible state of the system. From the vantage point of free energy, the exploration of phase transition conditions becomes accessible through the scrutiny of the emergence and stability of distinct phases. Furthermore, the characterization and analysis of the free energy landscape are profoundly influenced by its relationship with the order parameters. Taking the radius of a black hole’s event horizon as an exemplary order parameter, one can articulate the free energy landscape. Consequently, the phase transition of a black hole can be interpreted as a stochastic thermal fluctuation of said order parameter, offering a novel perspective on the underlying physics 21 . In the following, since we want to study the phase transition from the perspective of the free energy landscape and in a different structure under specific thermal fluctuations, we will introduce different forms of this energy that can help us in this way.
II.2.1 Gibbs Free Energy Landscape
@@
The terminology ’on-shell’ and ’off-shell’ delineates the aforementioned temperature discrepancy. ’Off-shell’ encompasses all transitions occurring at temperature ”T”, which, despite conforming to the laws of thermodynamics, do not necessarily constitute solutions to Einstein’s field equations 21 . It is imperative to recognize that solely the extremal points on the $G_{L}$ signify the authentic black hole phases that comply with the Einstein field equations. Furthermore, within this context, the local maxima and minima are indicative of the unstable and locally stable black hole phases, respectively 60 .
II.2.2 Landau free energy
The Landau free energy constitutes another showy path for scrutinizing the phase transition dynamics within the system’s free energy landscape. While using this method assumes heightened significance in the context of second-order phase transitions- wherein the bifurcation of the Landau functional’s local minimum epitomizes the thermodynamic system’s second-order phase transition- it is also pertinent to first-order phase transitions. Specifically, the transmutation of the global minimum is emblematic of a first-order phase transition, underscoring the multifaceted nature of the Landau-free energy in characterizing phase behavior. This function can be written as follows 45 ,
@@ II.2.2 Landau free energy
II.2.3 Thermal potential
Another form of landscape energy that can be used to investigate phase transition based on thermal fluctuation is the model that was introduced in 47 under the title of thermal potential.
@@ II.2.3 Thermal potential
III Non-linear magnetic-charged AdS black hole surrounded
@@
In this article, we went to black holes that somehow interact with the structure of matter and dark energy, so that we can study the effect of the presence of parameters caused by the dark structure on the first-order phase transition. So the metric of (NLM-C-Q-PFD) black holeis as follows 51 ,
@@
For the main quantities of this model i.e. mass $M$, Hawking temperature $T_{H}$, entropy $S$, pressure $P$ and volume $V$ we will have (note that as long as there is no specific definition of the radius, it will be $r=r_{H}$ in calculations) 51 ,
@@
III.0.1 Behavioral comparison of landscape free energy
Now according to Eqs. (II.1), (II.1), (II.1) and (II.1) for Gibbs and Landau free energy function and also the thermal potential of this model, we will have [^1] Note that in the following relationships, due to the number of parameters, the beginning of the equations is shown with a bold symbol only for better clarity.,